# Fluctuation-dissipation and virtual processes in interacting phonon systems

### v1

@@ Abstract
At the three-phonon level, the impact of the fluctuation-dissipation condition is particularly striking. While enforcing strict energy conservation allows us to reproduce theoretical results from the literature [24, 25, 15], incorporating self-consistent linewidths consistently leads to lower thermal conductivity values, which are in closer agreement with experimental measurements [23, 21, 22]. This reduction can be attributed to a significant decrease in the thermal conductivity contribution from phonons with energies around $20$ meV. Under strict energy conservation, these phonons are forbidden from interacting with optical modes, but the convolution in the self-consistent approach relaxes this restriction, thereby expanding the scattering phase space. This behavior is analogous to what occurs when four-phonon processes are included: the increased scattering phase space results in higher linewidths for these phonons and a corresponding reduction in their thermal conductivity contribution. Our findings challenge the prevailing view that four-phonon interactions are essential to explain the thermal conductivity in large phonon band gap materials like boron arsenide. While these interactions are not negligible, their importance appears to be overestimated. Instead, remaining at the three-phonon level while enforcing the fluctuation-dissipation condition provides a reasonably accurate prediction of thermal transport in this material, consistent with its low anharmonicity.
Although the fluctuation-dissipation condition is universally applicable, its impact on different materials varies, and requires quantitative evaluation. In the SM [7], we demonstrate that the impact of self-consistency is negligible for silicon — a highly harmonic system with low linewidths and no phonon band gap — whereas it is significant for silver iodide, which exhibits strong anharmonicity. Ultimately, we anticipate that self-consistency will substantially alter the phonon dynamics in systems with large mass differences or pronounced anharmonicity. For example, hydrogen-rich materials, which feature both characteristics, are promising candidates for further exploration of the fluctuation-dissipation condition’s role in phonon interactions.

### v2

@@ Abstract
At the three-phonon level, the impact of the fluctuation-dissipation condition is particularly striking. While enforcing strict energy conservation allows us to reproduce theoretical results from the literature [24, 25, 15], incorporating self-consistent linewidths consistently leads to lower thermal conductivity values, which are in closer agreement with experimental measurements [23, 21, 22]. This reduction can be attributed to a significant decrease in the thermal conductivity contribution from phonons with energies around $20$ meV. Under strict energy conservation, these phonons are forbidden from interacting with optical modes, but the convolution in the self-consistent approach relaxes this restriction, thereby expanding the scattering phase space. This behavior is analogous to what occurs when four-phonon processes are included: the increased scattering phase space results in higher linewidths for these phonons and a corresponding reduction in their thermal conductivity contribution. Our findings challenge the prevailing view that four-phonon interactions are essential to explain the thermal conductivity in large phonon band gap materials like boron arsenide. While these interactions are not negligible, their importance appears to have been overestimated in previous literature. Remaining at the three-phonon level while enforcing the fluctuation-dissipation condition provides a reasonably accurate prediction of thermal transport in this material, consistent with its low anharmonicity. However, this accuracy should not be seen as an indication that the self-consistent determination of linewidth could avoid the need for four-phonon interactions. Instead, the fluctuation-dissipation condition should be seen as a new, superior framework to evaluate all scattering mechanisms. Its importance in the context of BAs stems from the constrained 3-phonon scattering phase-space produced by strict energy conservation. The FD approach removes this constraint, and shows the reproduction of the experimental thermal conductivity of BAs using energy conservation and the 4-phonon channel is actually a coincidence: the EC 3-phonon $\kappa$ is too high, and the EC 4-phonon correction is too large, but happens to land near the experimental values.
The fluctuation-dissipation condition is universally applicable, but its impact on different materials varies, and requires quantitative evaluation. In the SM [7], we show the impact of the self-consistency for silicon, a highly harmonic system with low linewidths and no phonon band gap, and for silver iodide, which exhibits strong anharmonicity. In the former case, the difference in using the fluctuation-dissipation condition is negligible, both with three and four-phonon interactions. In AgI however, its impact is more pronounced. Including only three-phonon scattering, the self-consistency reduces the thermal conductivity by 15% to 25% between 100 to 300 K. When four phonon scattering is included, the impact of the fluctuation-dissipation condition is smaller, between 5% and 20 % in the same range of temperature, which is still significant for low thermal conductivity materials. Ultimately, the fluctuation-dissipation condition is a universal principle that applies to all materials, yet its impact on phonon dynamics appears to be highly system-dependent. Our findings highlight the need for broader investigations to determine how this condition interacts with anharmonicity and the available scattering phase space. In more complex or strongly interacting systems, additional factors, such as possible deviations from Markovian behavior, may also influence the outcome. Such studies will be essential for establishing when and how fluctuation-dissipation-consistent treatments are necessary for accurate modeling of lattice dynamics.